# Gradient flow for deep equilibrium single-index models

### v1

@@ 1 Introduction
The deep equilibrium model (DEQ) framework introduced in [48] has recently emerged as a powerful paradigm for training neural networks that has enjoyed practical success across many modern machine learning tasks, including image generation [66, 69], inverse problems [58], graph neural networks [51, 65], classification and semantic segmentation [50], representation learning [59], and large language modeling [48, 74]. DEQs have been shown to have several advantages over traditional deep neural networks, including lower memory consumption during training [57, 63], the ability to flexibly adapt to different computational budgets at inference time [71, 72, 74], and better representational power [73, 75].
Instead of learning an explicit relation $y_{i}=f(x_{i})$ from samples $\{(x_{i},y_{i})\}_{i=1}^{n}$, DEQs attempt to learn an implicit relation $y_{i}=g(x_{i},y_{i})$ over a parameterized class of functions $\{g_{\theta}\}_{\theta\in\Theta}$. Given an input $x$, the fixed point $y_{\theta}(x)$ of $g_{\theta}(x,\,\cdot)$ can often be interpreted as the limiting point of the sequence $y^{(t+1)}=g_{\theta}(x,y^{(t)})$ for any initial injection $y^{(0)}$; hence, we may view DEQs as a single layer which implicitly defines infinitely deep neural networks with input injection and weight tying across layers. In fact, [48] showed that there is no advantage to stacking single-layer DEQs. The training of DEQs is typically performed using the implicit function theorem, and inference is performed using a root-finding algorithm to solve for the fixed point $y_{\theta}(x)$.
Despite their practical success, theoretically understanding the gradient descent dynamics for training DEQs remains an area of active research [60, 70, 73]. In this work, we initiate a rigorous study of the gradient descent dynamics for DEQs in the simple setting of linear models (where the target is $f(x)=\xi^{\top}x$) and single-index models (where the target is $f(x)=\sigma(\xi^{\top}x)$ for some activation function $\sigma:\mathbb{R}\to\mathbb{R}$). We concern ourselves with the following questions:
@@ 2. Q2:
Using techniques from dynamical systems theory and previously developed tools for analyzing gradient flow for explicit single-neuron neural networks (e.g., [54]), we provide a full answer to Q1 and a partial answer to Q2 in the settings of well-specified linear and single-index models. Similarly to [54], we work directly with the population risk and focus on the gradient flow and gradient descent dynamics for minimizing this risk, leaving the analysis of empirical risk minimization to future work. Our work makes similar assumptions to prior work on gradient flow for single-index models, but notably, our final rates depend on the amount of nonlinearity present in the activation function. By analyzing both linear and nonlinear single-index models, we are therefore able to highlight the differences in the training dynamics of DEQs that arise due to the presence of nonlinearity.
@@ 1.2 Related work
DEQs were introduced in [48] as a framework for training infinitely deep neural networks with weight tying across layers and input injection. Since then, DEQs have been successfully applied to image generation [66, 69], inverse problems [58], graph neural networks [51, 65], classification and semantic segmentation [50], representation learning [59], and large language modeling [48, 74]. [76] recently suggest that it can be computationally beneficial to create an algebraically reversible fixed-point iteration, so that we can directly backpropagate through the fixed point iteration instead of approximating the gradient by the inverse problem in (1); however, their approach turns the implicit model into an explicit one, so it is not covered by our theory.
@@
There is a large body of work establishing that gradient descent converges for non-convex problems which satisfy a strict saddle property, meaning that all suboptimal saddle points admit a direction of easy escape [42, 43, 44]. For example, this phenomenon has been heavily studied for the phase retrieval problem (with $\sigma(z)=z^{2}$), where gradient descent is known to converge exponentially fast with appropriate spectral initialization [41, 47]. Gradient flow for explicit single-index models has been extensively studied in the literature [45, 49, 54], where typical assumptions include monotonicity of the activation function (which is assumed to have a strong gradient near zero) and that the data is sufficiently spread out near zero.
@@
[60] studies the optimization landscape of implicit models, showing that gradient descent converges at a linear rate to the global optimum for overparameterized linear implicit models. However, their results consider a stylized setting including a softmax nonlinearity on the weights in order to ensure well-posedness of the implicit model during training; we make no such modifications. In a similar vein, [70] and [64] show that the loss landscape for DEQs is benign in the heavily overparameterized regime where the width of the implicit layer is much larger than the number of samples. [68] shows that infinitely wide implicit models can be modeled by Gaussian processes, and [67] shows that there is a deterministic neural tangent kernel (NTK) for implicit models under mild conditions and show how to find it explicitly by root-finding.
Closest to our work, [73] analyzes the gradient descent dynamics for training diagonal linear DEQs and shows global convergence of gradient descent in the overparameterized regime. However, they analyze a different parameterization $y_{\theta}(x)=g_{\theta}(x,y_{\theta}(x))=\theta_{1}^{\top}x+\theta_{2}^{\top}y_{\theta}(x)\in\mathbb{R}^{d}$ and only assume access to samples—they do not assume that the ground truth is linear. As a result, they do not obtain the conservation law that we derive in this work or our explicit characterization of the limiting parameters for linear DEQs.
@@
There is a long line of work studying well-posedness of the Jacobian inverse in (1). For example, [53], [52], and [56] study sufficient conditions on the form of the neural network $g_{\theta}$ to ensure that the fixed-point problem is well-posed. [55] point out that regularizing the training of implicit models can prevent overfitting and training instabilities arising from ill-conditioned Jacobians. [62] show that initializing implicit models using orthogonal or symmetric matrices can help ensure that the Jacobian inverse is well-defined.
@@ 2 Background
For the backward pass, we would like to minimize $L(y_{\theta}(x))$ for some differentiable risk function $L$, by gradient descent. Given a loss function $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, we may define the population risk as
@@ 3.1.1 Gradient flow
One may wonder whether the assumption that $\mathbb{E}[XX^{\top}]\succ 0$ is necessary for this result; indeed, simulations show that if $\mathbb{E}[XX^{\top}]$ is rank-deficient, then it is still possible for the gradient flow to be well-defined for all time and converge. However, relaxing this assumption is nontrivial even for learning a single neuron with gradient flow [61], so we leave this to future work.
@@
$R(\theta(t),\xi)=\left(\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right)^{\top}\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right)\geq\lambda_{\mathrm{min}}(\mathbb{E}[XX^{\top}])\,\left\lVert\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right\rVert_{2}^{2}\to\infty.$
@@
so $V$ is a Lyapunov function for the gradient flow. Finally, the set of points where $\frac{d}{dt}V(\theta(t))=0$ is precisely the set of points $S=\{(\theta_{1},\theta_{2})\in\Omega:\theta_{1}=\xi(1-\theta_{2})\}$ by (3). By LaSalle’s invariance principle [39, Theorem 4.4], we conclude that the gradient flow converges to a point in $S$ as $t\to\infty$, which must be a global minimizer of the population risk since $\theta_{1}=\xi(1-\theta_{2})$. We can solve for this set by parameterizing $\theta=(\alpha\xi,\,1-\alpha)$ for $\alpha>0$:
@@
Finally, we can use this result to show exponentially fast convergence by Grönwall’s inequality, because convergence of the parameters implies that $\lvert 1-\theta_{2}(t)\rvert$ remains bounded away from zero.
@@
$=\frac{-2\,\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(t)}{(1-\theta_{2}(t))^{2}}-\frac{\xi}{1-\theta_{2}(t)}\right)\,(1-\theta_{2}(t))-2\,\theta_{1}(t)\,\frac{\theta_{1}(t)^{\top}}{(1-\theta_{2}(t))^{2}}\,\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right)}{(1-\theta_{2}(t))^{2}}$
@@ 3.1.2 Gradient descent
We can also analyze the gradient descent dynamics for fitting the linear deep equilibrium model (2). First, we recall the well-known descent lemma, which is a consequence of Theorem 2.1.5 of [46].
@@
$w(s+1)=\left\lVert\theta(s+1)-\begin{pmatrix}0\\ 1\end{pmatrix}\right\rVert_{2}^{2}$ $=\left\lVert\theta(s)-\eta\,\nabla_{\theta}\,R(\theta(s),\xi)-\begin{pmatrix}0\\ 1\end{pmatrix}\right\rVert_{2}^{2}$
@@
$\left(\theta(s)-\begin{pmatrix}0\\ 1\end{pmatrix}\right)^{\top}\,\nabla_{\theta}\,R(\theta(s),\xi)=2\,\begin{pmatrix}\theta_{1}(s)\\ \theta_{2}(s)-1\end{pmatrix}^{\top}\begin{pmatrix}\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(s)}{(1-\theta_{2}(s))^{2}}-\frac{\xi}{1-\theta_{2}(s)}\right)\\ \frac{\theta_{1}(s)^{\top}}{(1-\theta_{2}(s))^{2}}\,\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(s)}{1-\theta_{2}(s)}-\xi\right)\end{pmatrix}=0,$
@@
$\lVert\nabla_{\theta}\,R(\theta(t),\xi)\rVert_{2}^{2}$ $=2\,\left\lVert\frac{1}{1-\theta_{2}(t)}\begin{pmatrix}\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\\ \varphi(t)^{\top}\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\end{pmatrix}\right\rVert_{2}^{2}$
$=2\left(\left\lVert\frac{1}{1-\theta_{2}(t)}\,\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\right\rVert_{2}^{2}+\left\lVert\frac{\varphi(t)^{\top}}{1-\theta_{2}(t)}\,\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\right\rVert_{2}^{2}\right)$
$\geq\frac{2}{\alpha^{2}}\,\lambda_{\mathrm{min}}(\mathbb{E}[XX^{\top}])\,\lVert\varphi(t)-\xi\rVert_{2}^{2}.$
@@ 3.2.1 Gradient flow
    The data satisfies:
    • (i)
      $X\in L^{2}(\mathbb{P})$.
    • (ii)
      $\mathbb{E}[\lVert X\rVert_{2}^{2}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}]\geq\beta$ for some constants $\alpha,\beta\in(0,\infty)$.
@@
The assumptions on the data distribution are quite mild; for example, if $X\in L^{2}(\mathbb{P})$ is not almost surely equal to zero, then we can always find constants $\alpha,\beta$ satisfying the assumptions. The assumptions on the activation function rule out functions such as the linear activation $\sigma(z)=z$ (due to the nonlinearity condition), the step function $\sigma(z)=\mathbf{1}_{z\geq 0}$ (due to differentiability), and the ReLU activation $\sigma(z)=\max\{0,z\}$ (due to differentiability and the lower bound on the derivative near zero).
@@
Our proof technique is inspired by [54], and the goal is to apply Grönwall’s inequality to $r(t)\coloneq\lVert\theta(t)-(\xi,0)\rVert_{2}^{2}$. Throughout the proof, we will use the notation $\omega_{t}(z)\coloneq\theta_{1}(t)^{\top}z+\theta_{2}(t)\,y_{\theta(t)}(z)$. First, we compute
@@
$\begin{split}r^{\prime}(t)&\leq-\frac{2\gamma^{2}}{1+L\delta_{2}}\,\mathbb{E}\left[\left((\theta_{1}(t)-\xi)^{\top}X+\theta_{2}(t)\,y_{\theta(t)}(X)\right)^{2}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\\ &\leq-\frac{2\gamma^{2}}{1+L\delta_{2}}\,\mathbb{E}\left[\begin{pmatrix}\theta_{1}(t)-\xi\\ \theta_{2}(t)\end{pmatrix}^{\top}\begin{pmatrix}XX^{\top}&y_{\theta(t)}(X)\,X\\ y_{\theta(t)}(X)\,X^{\top}&y_{\theta(t)}(X)^{2}\end{pmatrix}\begin{pmatrix}\theta_{1}(t)-\xi\\ \theta_{2}(t)\end{pmatrix}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\\ &\leq-\frac{2\gamma^{2}}{1+L\delta_{2}}\,\lambda_{\mathrm{min}}\left(\mathbb{E}\left[\begin{pmatrix}XX^{\top}&y_{\theta(t)}(X)\,X\\ y_{\theta(t)}(X)\,X^{\top}&y_{\theta(t)}(X)^{2}\end{pmatrix}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\right)\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}.\end{split}$ (17)
@@
$\lambda_{\mathrm{min}}\left(\mathbb{E}\left[\begin{pmatrix}XX^{\top}&y_{\theta(t)}(X)\,X\\ y_{\theta(t)}(X)\,X^{\top}&y_{\theta(t)}(X)^{2}\end{pmatrix}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\right)=\inf_{u\in S^{d}}\;\mathbb{E}\left[\left(u^{\top}\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right)^{2}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right].$
@@
Since $X\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$ is not almost surely zero by our assumption on the data distribution, this expectation is strictly positive unless $y_{\theta(t)}(X)\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$ is almost surely equal to some linear function of $X\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$. Supposing for a contradiction that $y_{\theta(t)}(X)\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}=\kappa_{2}^{\top}X\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$ almost surely for some $\kappa_{2}\in\mathbb{R}^{d}$, we would have
$\kappa_{2}^{\top}X\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$ $=y_{\theta(t)}(X)\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$
$=\sigma(\theta_{1}(t)^{\top}X+\theta_{2}(t)\,y_{\theta(t)}(X))\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$
$=\sigma((\theta_{1}(t)+\theta_{2}(t)\,\kappa_{2})^{\top}X)\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha},$
contradicting our nonlinearity assumption on $\sigma$ if we pick $\kappa_{1}=\theta_{1}(t)+\theta_{2}(t)\,\kappa_{2}$. Hence, the expectation is strictly positive for any fixed $\theta(t)\in\Theta$. The dominated convergence theorem again implies that the expectation is continuous in $\theta$ over the compact set $\Theta$, so the extreme value theorem implies that there exists some constant $\rho>0$ such that
@@
Suppose that the activation function $\sigma\in C^{1}(\mathbb{R})$ is strictly increasing with $\lVert\sigma\rVert_{\mathrm{Lip}}=L<\infty$. Furthermore, suppose that $\sigma$ is not linear in any neighborhood of zero. Then, as long as $X$ is a continuous random variable which is not almost surely equal to zero and we initialize $\theta(0)\in B((\xi,0),1/L)$, the gradient flow for (10) converges exponentially fast to the target parameter $(\xi,0)$:

### v2

# Gradient descent for deep equilibrium single-index models

@@ 1 Introduction
The deep equilibrium model (DEQ) framework introduced in [undefau] has recently emerged as a powerful paradigm for training neural networks that has enjoyed practical success across many modern machine learning tasks, including image generation [undefaam, undefaap], inverse problems [undefaae], graph neural networks [undefax, undefaal], classification and semantic segmentation [undefaw], representation learning [undefaaf], and large language modeling [undefau, undefaau]. DEQs have been shown to have several advantages over traditional deep neural networks, including lower memory consumption during training [undefaad, undefaaj], the ability to flexibly adapt to different computational budgets at inference time [undefaar, undefaas, undefaau], and better representational power [undefaat, undefaav].
Instead of learning an explicit relation $y_{i}=f(x_{i})$ from samples $\{(x_{i},y_{i})\}_{i=1}^{n}$, DEQs attempt to learn an implicit relation $y_{i}=g(x_{i},y_{i})$ over a parameterized class of functions $\{g_{\theta}\}_{\theta\in\Theta}$. Given an input $x$, the fixed point $y_{\theta}(x)$ of $g_{\theta}(x,\,\cdot)$ can often be interpreted as the limiting point of the sequence $y^{(t+1)}=g_{\theta}(x,y^{(t)})$ for any initial injection $y^{(0)}$; hence, we may view DEQs as a single layer which implicitly defines infinitely deep neural networks with input injection and weight tying across layers. In fact, [undefau] showed that there is no advantage to stacking single-layer DEQs. The training of DEQs is typically performed using the implicit function theorem, and inference is performed using a root-finding algorithm to solve for the fixed point $y_{\theta}(x)$.
Despite their practical success, theoretically understanding the gradient descent dynamics for training DEQs remains an area of active research [undefaag, undefaaq, undefaat]. In this work, we initiate a rigorous study of the gradient descent dynamics for DEQs in the simple setting of linear models (where the target is $f(x)=\xi^{\top}x$) and single-index models (where the target is $f(x)=\sigma(\xi^{\top}x)$ for some activation function $\sigma:\mathbb{R}\to\mathbb{R}$). We concern ourselves with the following questions:
@@ 2. Q2:
Using techniques from dynamical systems theory and previously developed tools for analyzing gradient flow for explicit single-neuron neural networks (e.g., [undefaaa]), we provide a full answer to Q1 and a partial answer to Q2 in the settings of well-specified linear and single-index models. Similarly to [undefaaa], we work directly with the population risk and focus on the gradient flow and gradient descent dynamics for minimizing this risk, leaving the analysis of empirical risk minimization to future work. Our work makes similar assumptions to prior work on gradient flow for single-index models, but notably, our final rates depend on the amount of nonlinearity present in the activation function. By analyzing both linear and nonlinear single-index models, we are therefore able to highlight the differences in the training dynamics of DEQs that arise due to the presence of nonlinearity.
@@ 1.2 Related work
DEQs were introduced in [undefau] as a framework for training infinitely deep neural networks with weight tying across layers and input injection. Since then, DEQs have been successfully applied to image generation [undefaam, undefaap], inverse problems [undefaae], graph neural networks [undefax, undefaal], classification and semantic segmentation [undefaw], representation learning [undefaaf], and large language modeling [undefau, undefaau]. [undefaaw] recently suggest that it can be computationally beneficial to create an algebraically reversible fixed-point iteration, so that we can directly backpropagate through the fixed point iteration instead of approximating the gradient by the inverse problem in (1); however, their approach turns the implicit model into an explicit one, so it is not covered by our theory.
@@
There is a large body of work establishing that gradient descent converges for non-convex problems which satisfy a strict saddle property, meaning that all suboptimal saddle points admit a direction of easy escape [undefao, undefap, undefaq]. For example, this phenomenon has been heavily studied for the phase retrieval problem (with $\sigma(z)=z^{2}$), where gradient descent is known to converge exponentially fast with appropriate spectral initialization [undefan, undefat]. Gradient flow for explicit single-index models has been extensively studied in the literature [undefar, undefav, undefaaa], where typical assumptions include monotonicity of the activation function (which is assumed to have a strong gradient near zero) and that the data is sufficiently spread out near zero.
@@
[undefaag] studies the optimization landscape of implicit models, showing that gradient descent converges at a linear rate to the global optimum for overparameterized linear implicit models. However, their results consider a stylized setting including a softmax nonlinearity on the weights in order to ensure well-posedness of the implicit model during training; we make no such modifications. In a similar vein, [undefaaq] and [undefaak] show that the loss landscape for DEQs is benign in the heavily overparameterized regime where the width of the implicit layer is much larger than the number of samples. [undefaao] shows that infinitely wide implicit models can be modeled by Gaussian processes, and [undefaan] shows that there is a deterministic neural tangent kernel (NTK) for implicit models under mild conditions and show how to find it explicitly by root-finding.
Closest to our work, [undefaat] analyzes the gradient descent dynamics for training diagonal linear DEQs and shows global convergence of gradient descent in the overparameterized regime. However, they analyze a different parameterization $y_{\theta}(x)=g_{\theta}(x,y_{\theta}(x))=\theta_{1}^{\top}x+\theta_{2}^{\top}y_{\theta}(x)\in\mathbb{R}^{d}$ and only assume access to samples—they do not assume that the ground truth is linear. As a result, they do not obtain the conservation law that we derive in this work or our explicit characterization of the limiting parameters for linear DEQs.
@@
There is a long line of work studying well-posedness of the Jacobian inverse in (1). For example, [undefaz], [undefay], and [undefaac] study sufficient conditions on the form of the neural network $g_{\theta}$ to ensure that the fixed-point problem is well-posed. [undefaab] point out that regularizing the training of implicit models can prevent overfitting and training instabilities arising from ill-conditioned Jacobians. [undefaai] show that initializing implicit models using orthogonal or symmetric matrices can help ensure that the Jacobian inverse is well-defined.
@@ 2 Background
For the backward pass, we would like to minimize the expected loss by gradient descent. Given a loss function $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, we may define the population risk as
@@ 3.1.1 Gradient flow
One may wonder whether the assumption that $\mathbb{E}[XX^{\top}]\succ 0$ is necessary for this result; indeed, simulations show that if $\mathbb{E}[XX^{\top}]$ is rank-deficient, then it is still possible for the gradient flow to be well-defined for all time and converge. However, relaxing this assumption is nontrivial even for learning a single neuron with gradient flow [undefaah], so we leave this to future work.
@@
$R(\theta(t),\xi)$ $=\left(\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right)^{\top}\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right)$
$\geq\lambda_{\mathrm{min}}(\mathbb{E}[XX^{\top}])\,\left\lVert\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right\rVert_{2}^{2}$
$\to\infty.$
@@
so $V$ is a Lyapunov function for the gradient flow. Finally, the set of points where $\frac{d}{dt}V(\theta(t))=0$ is precisely the set of points $S=\{(\theta_{1},\theta_{2})\in\Omega:\theta_{1}=\xi(1-\theta_{2})\}$ by (3). By LaSalle’s invariance principle [undefal, Theorem 4.4], we conclude that the gradient flow converges to a point in $S$ as $t\to\infty$, which must be a global minimizer of the population risk since $\theta_{1}=\xi(1-\theta_{2})$. We can solve for this set by parameterizing $\theta=(\alpha\xi,\,1-\alpha)$ for $\alpha>0$:
@@
Finally, we can use this result to show exponentially fast convergence using Grönwall’s inequality, because convergence of the parameters implies that $\lvert 1-\theta_{2}(t)\rvert$ remains bounded away from zero.
@@
$=-\frac{2\,\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(t)}{(1-\theta_{2}(t))^{2}}-\frac{\xi}{1-\theta_{2}(t)}\right)\,(1-\theta_{2}(t))}{(1-\theta_{2}(t))^{2}}$
$\qquad-\frac{2\,\theta_{1}(t)\,\frac{\theta_{1}(t)^{\top}}{(1-\theta_{2}(t))^{2}}\,\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(t)}{1-\theta_{2}(t)}-\xi\right)}{(1-\theta_{2}(t))^{2}}$
@@ 3.1.2 Gradient descent
We can also analyze the gradient descent dynamics for fitting the linear deep equilibrium model (2). First, we recall the well-known descent lemma, which is a consequence of Theorem 2.1.5 of [undefas].
@@
$w(s+1)$ $=\left\lVert\theta(s+1)-\begin{pmatrix}0\\ 1\end{pmatrix}\right\rVert_{2}^{2}$
$=\left\lVert\theta(s)-\eta\,\nabla_{\theta}\,R(\theta(s),\xi)-\begin{pmatrix}0\\ 1\end{pmatrix}\right\rVert_{2}^{2}$
@@
$\left(\theta(s)-\begin{pmatrix}0\\ 1\end{pmatrix}\right)^{\top}\,\nabla_{\theta}\,R(\theta(s),\xi)$
$\quad=2\,\begin{pmatrix}\theta_{1}(s)\\ \theta_{2}(s)-1\end{pmatrix}^{\top}\begin{pmatrix}\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(s)}{(1-\theta_{2}(s))^{2}}-\frac{\xi}{1-\theta_{2}(s)}\right)\\ \frac{\theta_{1}(s)^{\top}}{(1-\theta_{2}(s))^{2}}\,\mathbb{E}[XX^{\top}]\left(\frac{\theta_{1}(s)}{1-\theta_{2}(s)}-\xi\right)\end{pmatrix}$
$\quad=0,$
@@
$\lVert\nabla_{\theta}\,R(\theta(t),\xi)\rVert_{2}^{2}$
$\quad=2\,\left\lVert\frac{1}{1-\theta_{2}(t)}\begin{pmatrix}\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\\ \varphi(t)^{\top}\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\end{pmatrix}\right\rVert_{2}^{2}$
$\quad=2\left(\left\lVert\frac{1}{1-\theta_{2}(t)}\,\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\right\rVert_{2}^{2}+\left\lVert\frac{\varphi(t)^{\top}}{1-\theta_{2}(t)}\,\mathbb{E}[XX^{\top}]\,(\varphi(t)-\xi)\right\rVert_{2}^{2}\right)$
$\quad\geq\frac{2}{\alpha^{2}}\,\lambda_{\mathrm{min}}(\mathbb{E}[XX^{\top}])\,\lVert\varphi(t)-\xi\rVert_{2}^{2}.$
@@ 3.2.1 Gradient flow
    The data satisfies $X\in L^{2}(\mathbb{P})$ and $\mathbb{E}[XX^{\top}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}]\succ 0$ for some $\alpha>0$.
@@
First, we only assume that the covariance of the data is nondegenerate. The assumptions on the activation function rule out functions such as the linear activation $\sigma(z)=z$ (due to the nonlinearity condition), the step function $\sigma(z)=\mathbf{1}_{z\geq 0}$ (due to differentiability), and the ReLU activation $\sigma(z)=\max\{0,z\}$ (due to differentiability and the lower bound on the derivative near zero).
@@
Our proof technique is inspired by [undefaaa], and the goal is to apply Grönwall’s inequality to $r(t)\coloneq\lVert\theta(t)-(\xi,0)\rVert_{2}^{2}$. Throughout the proof, we will use the notation $\omega_{t}(z)\coloneq\theta_{1}(t)^{\top}z+\theta_{2}(t)\,y_{\theta(t)}(z)$. First, we compute
@@
$\begin{split}&r^{\prime}(t)\\ &\quad\leq-\frac{2\gamma^{2}}{1+L\delta_{2}}\,\mathbb{E}\left[\left((\theta_{1}(t)-\xi)^{\top}X+\theta_{2}(t)\,y_{\theta(t)}(X)\right)^{2}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\\ &\quad=-\frac{2\gamma^{2}}{1+L\delta_{2}}\,\mathbb{E}\left[\begin{pmatrix}\theta_{1}(t)-\xi\\ \theta_{2}(t)\end{pmatrix}^{\top}\begin{pmatrix}XX^{\top}&y_{\theta(t)}(X)\,X\\ y_{\theta(t)}(X)\,X^{\top}&y_{\theta(t)}(X)^{2}\end{pmatrix}\begin{pmatrix}\theta_{1}(t)-\xi\\ \theta_{2}(t)\end{pmatrix}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\\ &\quad\leq-\frac{2\gamma^{2}}{1+L\delta_{2}}\,\lambda_{\mathrm{min}}\left(\mathbb{E}\left[\begin{pmatrix}XX^{\top}&y_{\theta(t)}(X)\,X\\ y_{\theta(t)}(X)\,X^{\top}&y_{\theta(t)}(X)^{2}\end{pmatrix}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\right)\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}.\end{split}$ (17)
@@
$\lambda_{\mathrm{min}}\left(\mathbb{E}\left[\begin{pmatrix}XX^{\top}&y_{\theta(t)}(X)\,X\\ y_{\theta(t)}(X)\,X^{\top}&y_{\theta(t)}(X)^{2}\end{pmatrix}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]\right)$
$\quad=\inf_{u\in S^{d}}\;\mathbb{E}\left[\left(u^{\top}\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right)^{2}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right].$
@@
We now focus on the covariance matrix
$\Sigma=\begin{pmatrix}\Sigma_{XX}&\Sigma_{Xy}\\ \Sigma_{Xy}^{\top}&\Sigma_{yy}\end{pmatrix}\coloneq\mathbb{E}\left[\begin{pmatrix}XX^{\top}&y_{\theta(t)}(X)\,X\\ y_{\theta(t)}(X)\,X^{\top}&y_{\theta(t)}(X)^{2}\end{pmatrix}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right].$
Since we assumed that $\Sigma_{XX}\succ 0$, it follows that $\Sigma$ is singular if and only if the Schur complement $\Sigma_{yy}-\Sigma_{Xy}^{\top}\Sigma_{XX}^{-1}\Sigma_{Xy}$ is zero. Setting $\kappa_{2}\coloneq\Sigma_{XX}^{-1}\Sigma_{Xy}$, we explicitly calculate the mean squared error of using $\kappa_{2}^{\top}X$ to predict $y_{\theta(t)}(X)$:
$\mathbb{E}\left[(y_{\theta(t)}(X)-\kappa_{2}^{\top}X)^{2}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}\right]$
$\quad=\mathbb{E}[y_{\theta(t)}(X)^{2}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}]-2\kappa_{2}^{\top}\mathbb{E}[Xy_{\theta(t)}(X)\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}]+\kappa_{2}^{\top}\mathbb{E}[XX^{\top}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}]\kappa_{2}$
$\quad=\Sigma_{yy}-2\kappa_{2}^{\top}\Sigma_{Xy}+\kappa_{2}^{\top}\Sigma_{XX}\kappa_{2}$
$\quad=\Sigma_{yy}-2(\Sigma_{Xy}^{\top}\Sigma_{XX}^{-1})\Sigma_{Xy}+(\Sigma_{Xy}^{\top}\Sigma_{XX}^{-1})\Sigma_{XX}(\Sigma_{XX}^{-1}\Sigma_{Xy})$
$\quad=\Sigma_{yy}-2\Sigma_{Xy}^{\top}\Sigma_{XX}^{-1}\Sigma_{Xy}+\Sigma_{Xy}^{\top}\Sigma_{XX}^{-1}\Sigma_{Xy}$
$\quad=\Sigma_{yy}-\Sigma_{Xy}^{\top}\Sigma_{XX}^{-1}\Sigma_{Xy},$
which we recognize as the Schur complement. If the Schur complement is zero, then $(y_{\theta(t)}(X)-\kappa_{2}^{\top}X)^{2}\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}=0$ almost surely, implying $y_{\theta(t)}(X)=\kappa_{2}^{\top}X$ almost surely when $\lVert X\rVert_{2}\leq\alpha$. Substituting this into the fixed-point equation yields $\sigma((\theta_{1}(t)+\theta_{2}(t)\,\kappa_{2})^{\top}X)\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}=\kappa_{2}^{\top}X\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}$, which contradicts Assumption 3 if we set $\kappa_{1}=\theta_{1}(t)+\theta_{2}(t)\,\kappa_{2}$.
Hence, the minimum eigenvalue of the covariance is strictly positive for any fixed $\theta(t)\in\Theta$. The dominated convergence theorem again implies that the minimum eigenvalue is continuous in $\theta$ over the compact set $\Theta$, so the extreme value theorem implies that there exists some constant $\rho>0$ such that
@@
Suppose that the activation function $\sigma\in C^{1}(\mathbb{R})$ is strictly increasing with $\lVert\sigma\rVert_{\mathrm{Lip}}=L<\infty$. Furthermore, suppose that $\sigma$ is not linear in any neighborhood of zero. Then, as long as $X$ is a continuous random variable which has $\mathbb{E}[XX^{\top}]\succ 0$ and we initialize $\theta(0)\in B((\xi,0),1/L)$, the gradient flow for (10) converges exponentially fast to the target parameter $(\xi,0)$: